If a $10,000 par T-bill has a 4.25% discount quote and a 180-day maturity, what is the price of the T-bill to the nearest dollar?

Please show equations/work
Answer #1

Number of days to maturity = 180

Annual Discount rate = 4.25%

Effective Discount Rate for 180 days = 4.25 * (180/360) = 2.125%

Par Value = $10000

Hence, Price of T Bill = 10000* (1 - 0.02125) = $9787.50 or $9,788